OpenStudy (anonymous):

Solve using the one to one property of Logarithms. Please show steps. Ln(x+3)+Ln(x+1)=Ln8

OpenStudy (anonymous):

Ln((x+3)(x+1))=Ln8 (property used abv: Ln(a*b)=Lna+Lnb ) taking anti log (x+3)(x+1)=8... rest u cn solve....